# Discrete Optimal Transport and Voice Conversion

### v2

@@ 1 Introduction
In paper [1], the main deep learning approaches to VC are summarized. Most methods operate on spectrogram representations, often leveraging speaker embeddings or fundamental frequency (F0) information. Many of these techniques are inspired by image style transfer using generative adversarial networks (GANs), where a generator learns to produce outputs with the style of a target domain.
More recently, the neural optimal transport (NOT) framework was introduced (see [2] and references therein). NOT can be seen as a generalization of GANs: while GANs map noise to a target distribution, NOT explicitly learns a transformation between two data distributions. The application of NOT to VC was proposed in [3].
The emergence of vector-based audio representations, such as those produced by the wav2vec model [4], has opened new possibilities for VC. The HuBERT model [5] was used for controllable VC in article [6], while the WavLM model [7] was employed in work [8]. In the latter, each source vector was mapped to the average of its $k$ nearest neighbors in the target set. In paper [3], this was refined by selecting $k$ vectors based on the discrete optimal transport (OT) plan instead of simple kNN, and using flow matching to generalize to unseen data (see Section 2.2). In both works, the number of neighbors was fixed at $k=4$, with no ablation study performed.
@@
Additionally, we explore the ability of the discrete OT to work on more distinct unpaired domains. We apply the OT-based mapping to convert generated audio samples into the domain of real speech using the ASVspoof 2019 dataset. As a result, in the majority of cases, the AASIST model [9] misclassified the converted audio as bona fide speech, indicating the effectiveness of the proposed approach for domain adoptation (see Section 4.2) and introducing a new adversarial attack.
@@ 2.1 Discrete Optimal Transport
To compute the plan $\gamma,$ we use entropic OT with Sinkhorn algorithm described in [10, Ch. 4].
@@ 2.2 Continuous Optimal Transport
Under mild conditions, the solution of the Kantorovich OT problem follows from the solution of the Monge problem [11, Th. 5.30]. When paired examples are available (e.g., from discrete OT), flow matching method can be used to find an approximate continuous transport map $T,$ as in [12, Th. 4.2]. For neural network approximation of the transportation map see paper [2].
@@ 3.1 Interface
We use wav2vec audio representation, specifically the WavLM Large pretrained model [7]. This model encodes every $25$ ms of audio into a $1024$-dimensional vector embedding, with a hop size of $20$ ms. In addition to automatic speech recognition (ASR), WavLM was also trained for speaker identification, which allows it to preserve speaker identity in its embeddings.
@@ 3.4 Transportation Map
In KNN-VC approach [8], for each source embedding $x_{i},$ the target embeddings $y_{j}$ are sorted by decreasing cosine similarity. Denote these sorted vectors by $y_{j}^{knn(i)}.$ Then a $k$-nearest neighbors regression is applied:
@@
The approach used in paper [3] to obtain the training pairs for flow matching, we call OT-AVE. It averages over the top-$k$ target vectors,
@@ 4.1 Voice Conversion on LibriSpeech
We conduct our experiments using the LibriSpeech train-clean-100 dataset [13]. Following the protocol in [3], we select the first 40 speakers (ordered by speaker ID) and, for each, extract 10 random utterances and sort them by duration. We adopt an any-to-any conversion setup, in which each speaker is converted into the voice of the remaining 39 speakers.
@@
For each case, we perform an ablation study over different values of $k$. The Word Error Rate (WER)[14] and Mean Opinion Score (MOS)[15] results are shown in Fig. 1Fig. 3. In addition, for Case 3 (the full set of utterances), we report the Fréchet Audio Distance (FAD)[16] in Table 1, as it provides a more reliable estimate when sufficient embeddings are available.
@@
    Source duration $>$ 1 minute, target duration $<$ 1 minute (Fig. 4),
@@
    Source duration $<$ 1 minute, target duration $>$ 1 minute (Fig. 5).
Figure 2, Fig. 3, and Fig. 5 show the lowest WER, while the last two exhibit the highest MOS. These results align with prior findings in [8] and [3], suggesting that the duration of the target audio plays a crucial role in VC quality. Furthermore, the OT-BAR method consistently produces embedding distributions that are closer to the original target across most $k$ values.
@@ 4.2 Voice Conversion on ASVspoof
To test our method in a more challenging, domain-mismatched setting, we evaluate on the ASVspoof 2019 dataset [17]. We sort recordings by duration and select only those longer than 2.9 seconds, based on earlier observations that longer target utterances improve performance. We convert the first 1000 spoofed (fake) recordings into the last 1000 bona fide recordings using a one-to-one setup. The converted audio is then passed through the AASIST model for spoof detection. There are AASIST2[18] and AASIST3[19] models, but their code is not publicly available.
@@ 5.1 Datasets
For the experiments in Section 4.1 we used LibriSpeech Clean dataset available on Kaggle [20]. For the experiments in Section 4.2 we used ASVspoof 2019 dataset, also accessed via Kaggle [21].
@@ 5.2 Pipeline
We used the WavLM Large extracting embeddings from the sixth transformer layer and HiFi-GAN models from GitHub repository accompanying KNN-VC model [3].
The Sinkhorn algorithm, implemented in the POT library [22] (installed via pip install pot), was applied with a regularization parameter of $0.1$
The official AASIST model implementation [9] was used.
@@ 5.3 Metrics Evaluation
To compute WER, we used the Whisper speech recognition model [23] for transcription and the JiWER library (pip install jiwer) for comparison with reference transcriptions.
MOS was automatically computed using the code provided with the UTMOSv2 paper [24].
To calculate embeddings for FAD, we used the torchvggish 0.2 implementation (pip install -U torchvggish) of the VGGish model [25]. After downloading the model, we disabled postprocessing quantization and used the output from the last layer without activation. The embeddings were collected in 16-bit precision, converted to 32-bit, and used for means and covariance matrices calculation exactly in the same manner as in the official code for paper [26].
@@ 6 Conclusion
We also observed that the quality of converted speech, as measured by MOS and WER, is highly dependent on the duration of the target utterances. This supports prior findings [8, 3], and highlights the importance of sufficient target speaker data for achieving natural and intelligible outputs.
@@ 7 Acknowledgment
The first author thanks Mona Udasi for discussions related to the results in Section 4.2, Meiying Chen for pointing to the automatic MOS evaluation method [24], and Research Computing at the Rochester Institute of Technology for providing computational resources that supported this research.

### v3

@@ 1 Introduction
In paper [walczyna], the main deep learning approaches to VC are summarized. Most methods operate on spectrogram representations, often leveraging speaker embeddings or fundamental frequency (F0) information. Many of these techniques are inspired by image style transfer using generative adversarial networks (GANs), where a generator learns to produce outputs with the style of a target domain.
More recently, the neural optimal transport (NOT) framework was introduced (see [korotin2023neural] and references therein). NOT can be seen as a generalization of GANs: while GANs map noise to a target distribution, NOT explicitly learns a transformation between two data distributions. The application of NOT to VC was proposed in [VCOT].
The emergence of vector-based audio representations, such as those produced by the wav2vec model [wav2vec], has opened new possibilities for VC. The HuBERT model [hubert] was used for controllable VC in article [chen23r], while the WavLM model [wavlm] was employed in work [VCNN]. In the latter, each source vector was mapped to the average of its $k$ nearest neighbors in the target set. In paper [VCOT], this was refined by selecting $k$ vectors based on the discrete optimal transport (OT) plan instead of simple kNN, and using flow matching to generalize to unseen data (see Section 2.2). In both works, the number of neighbors was fixed at $k=4$, with no ablation study performed.
@@
Additionally, we explore the ability of the discrete OT to work on more distinct unpaired domains. We apply the OT-based mapping to convert generated audio samples into the domain of real speech using the ASVspoof 2019 dataset. As a result, in the majority of cases, the AASIST model [Jung2021AASIST] misclassified the converted audio as bona fide speech, indicating the effectiveness of the proposed approach for domain adoptation (see Section 4.2) and introducing a new adversarial attack.
@@ 2.1 Discrete Optimal Transport
To compute the plan $\gamma,$ we use entropic OT with Sinkhorn algorithm described in [ot:book, Ch. 4].
@@ 2.2 Continuous Optimal Transport
Under mild conditions, the solution of the Kantorovich OT problem follows from the solution of the Monge problem [villani:book, Th. 5.30]. When paired examples are available (e.g., from discrete OT), flow matching method can be used to find an approximate continuous transport map $T,$ as in [flowm, Th. 4.2]. For neural network approximation of the transportation map see paper [korotin2023neural].
@@ 3.1 Interface
We use wav2vec audio representation, specifically the WavLM Large pretrained model [wavlm]. This model encodes every $25$ ms of audio into a $1024$-dimensional vector embedding, with a hop size of $20$ ms. In addition to automatic speech recognition (ASR), WavLM was also trained for speaker identification, which allows it to preserve speaker identity in its embeddings.
@@ 3.4 Transportation Map
In KNN-VC approach [VCNN], for each source embedding $x_{i},$ the target embeddings $y_{j}$ are sorted by decreasing cosine similarity. Denote these sorted vectors by $y_{j}^{knn(i)}.$ Then a $k$-nearest neighbors regression is applied:
@@
The approach used in paper [VCOT] to obtain the training pairs for flow matching, we call OT-AVE. It averages over the top-$k$ target vectors,
@@ 4.1 Voice Conversion on LibriSpeech
We conduct our experiments using the LibriSpeech train-clean-100 dataset [panayotov2015librispeech]. Following the protocol in [VCOT], we select the first 40 speakers (ordered by speaker ID) and, for each, extract 10 random utterances and sort them by duration. We adopt an any-to-any conversion setup, in which each speaker is converted into the voice of the remaining 39 speakers.
@@
For each case, we perform an ablation study over different values of $k$. The Word Error Rate (WER)[wer] and Mean Opinion Score (MOS)[mos] results are shown in Fig. 1Fig. 3. In addition, for Case 3 (the full set of utterances), we report the Fréchet Audio Distance (FAD)[fad] in Table 1, as it provides a more reliable estimate when sufficient embeddings are available.
@@
    Source duration >> 1 minute, target duration << 1 minute (Fig. 4),
@@
    Source duration << 1 minute, target duration >> 1 minute (Fig. 5).
Figure 2, Fig. 3, and Fig. 5 show the lowest WER, while the last two exhibit the highest MOS. These results align with prior findings in [VCNN] and [VCOT], suggesting that the duration of the target audio plays a crucial role in VC quality. Furthermore, the OT-BAR method consistently produces embedding distributions that are closer to the original target across most $k$ values.
@@ 4.2 Voice Conversion on ASVspoof
To test our method in a more challenging, domain-mismatched setting, we evaluate on the ASVspoof 2019 dataset [wang2020asvspoof]. We sort recordings by duration and select only those longer than 2.9 seconds, based on earlier observations that longer target utterances improve performance. We convert the first 1000 spoofed (fake) recordings into the last 1000 bona fide recordings using a one-to-one setup. The converted audio is then passed through the AASIST model for spoof detection. There are AASIST2[aasist2] and AASIST3[aasist3] models, but their code is not publicly available.
@@ 5.1 Datasets
For the experiments in Section 4.1 we used LibriSpeech Clean dataset available on Kaggle [kaggleLS]. For the experiments in Section 4.2 we used ASVspoof 2019 dataset, also accessed via Kaggle [kaggleAS].
@@ 5.2 Pipeline
We used the WavLM Large extracting embeddings from the sixth transformer layer and HiFi-GAN models from GitHub repository accompanying KNN-VC model [VCOT].
The Sinkhorn algorithm, implemented in the POT library [pot] (installed via pip install pot), was applied with a regularization parameter of $0.1.$
The official AASIST model implementation [Jung2021AASIST] was used.
@@ 5.3 Metrics Evaluation
To compute WER, we used the Whisper speech recognition model [whisper] for transcription and the JiWER library (pip install jiwer) for comparison with reference transcriptions.
MOS was automatically computed using the code provided with the UTMOSv2 paper [utmosv2].
To calculate embeddings for FAD, we used the torchvggish 0.2 implementation (pip install -U torchvggish) of the VGGish model [vgg]. After downloading the model, we disabled postprocessing quantization and used the output from the last layer without activation. The embeddings were collected in 16-bit precision, converted to 32-bit, and used for means and covariance matrices calculation exactly in the same manner as in the official code for paper [fadtk].
@@ 6 Conclusion
We also observed that the quality of converted speech, as measured by MOS and WER, is highly dependent on the duration of the target utterances. This supports prior findings [VCNN, VCOT], and highlights the importance of sufficient target speaker data for achieving natural and intelligible outputs.
@@ 7 Acknowledgment
The first author thanks Mona Udasi for discussions related to the results in Section 4.2, Meiying Chen for pointing to the automatic MOS evaluation method [utmosv2], and Research Computing at the Rochester Institute of Technology for providing computational resources that supported this research.